OpenStudy (anonymous):

The curve y^2= 12x intersects the line 3y = 4x+ 6 at two points. Find the distance between the two points. Thanks for any help in advance.

OpenStudy (raden):

elimination and subtitution method is work

OpenStudy (raden):

from the equation of line : 3y = 4x+ 6 y = (4x+6)/3 ... (1) subtitute the value of y above into the equation y^2 = 12x, we get ((4x+6)/3)^2 = 12x (16x^2 + 48x + 36)/9 = 12x 16x^2 + 48x + 36 = 9 * 12x 16x^2 + 48x + 36 = 108x 16x^2 + 48x + 36 - 108x = 0 16x^2 -60x + 36 = 0 divide by 4 on both sides, simplied be 4x^2 - 15x + 9 = 0 now, can you find the solution for x from the equ above?

OpenStudy (anonymous):

(4x-3)(x-3) x=3/4, 3 Coordinates = (3/4, 3) and (3, 6) Distance= Distance = SquareRoot(2.25^2+3^2)=3.75 Thanks for your help RadEn!

OpenStudy (raden):

waw, you got it... im not yet get it, hehe.. you're welcome